OpenStudy (anonymous):

\[\int\limits {u \over 1+u} du\]

OpenStudy (anonymous):

take 1+u=v then u=v-1.. solve

OpenStudy (mimi_x3):

i think thats its long division..

sam (.sam.):

yes

OpenStudy (anonymous):

\[Original~ question~~~ \int\limits {e^x \over 1+e^x}dx--->x=Inu\]

sam (.sam.):

Once you do long div, you'll get \[\text{}=\int\limits \left(1-\frac{1}{x+1}\right) \, dx\]

sam (.sam.):

u=e^x+1and \[du=e^xd x\] \[\int\limits \frac{1}{u} \, du\]